Suppose a stock had an initial price of $88 per share, paid a dividend of $1.90 per share during the year, and had an ending share price of $73.50.

a. Compute the percentage total return. (A negative answer should be indicated by a minus sign. Do not round intermediate calculations and enter your answer as a percent rounded to 2 decimal places, e.g., 32.16.)
b. What was the dividend yield? (Do not round intermediate calculations and enter your answer as a percent rounded to 2 decimal places, e.g., 32.16.)
c. What was the capital gains yield? (A negative answer should be indicated by a minus sign. Do not round intermediate calculations and enter your answer as a percent rounded to 2 decimal places, e.g., 32.16.)
Answer #1

a) Percentage total return:-

=(ending price-intitial price+dividend)/initial price

=(73.5-88+1.9)/88

=-14.32%

b) Dividend yield

=1.9/88

=2.16

c) Capital yield;-

=73.5/88-1

=-16.48%